A fair six-sided dice is thrown and the scores are noted.

Event X: The total of the two scores is 4.

Even Y: The first score is 2 or 5.

a) Are events X and Y independent


1
Expert's answer
2021-07-06T16:53:12-0400

Solution:

We assume the die is thrown twice.

X = {(1,3),(3,1),(2,2)

Y = {(2,1),(2,2),(2,3),(2,4),(2,5),(2,6),(5,1),(5,2),(5,3),(5,4),(5,5),(5,6)}

Now, X "\cap" Y = {(2,2)} "\ne \phi"

And we know that if A and B are two independent events, then "P(A).P(B)=P(A\cap B)"

Here, "P(X)=\dfrac{3}{36}=\dfrac{1}{12},P(Y)=\dfrac{12}{36}=\dfrac{1}{3}"

And "P(X\cap Y)=\dfrac{1}{36}"

Clearly, "P(X).P(Y)=P(X\cap Y)=\dfrac{1}{36}"

So, events "X, \,Y" are independent.